An airplane is 12,000 ft horizontally from an airport. The slanted distance to the airport is 13,000 ft. What is the planes altitude?
Answer by checkley75(3666)

you have a right triangle with the hypotenuse=13,000, one leg=12,000 & the other leg=x.
using the pythagoreon theorum a^2+b^2=c^2 we get:
12,000^2+x^2=13,000^2 now we sove for x
144,000,000+x^2=169,000,000
x^2=169,000,000-144,000,000
x^2=25,000,000
x=sqrt25,000,000
x=5,000 ft. is the plane's altitude.